# DeContext as Defense: Safe Image Editing in Diffusion Transformers

### v1

@@ 1 Introduction
However, the very strength of in-context personalization also introduces serious privacy risks. Since they operate at inference time and require only a single image, adversaries can easily exploit publicly shared photos to clone identities, mimic artistic styles, or reproduce proprietary content without consent. Consequently, as shown in Figure LABEL:fig:results, the images users share online to express themselves may be weaponized to generate deepfakes, impersonation, and copyright violation, entirely bypassing conventional data protection mechanism.
@@ 2.1 Conditional Image Generation
Diffusion models have revolutionized controllable image synthesis [latentDM, SDXL, imagen, ediffi]. While early approaches rely on text prompts [GLIDE, DALLE2, cogview], recent advances enable diverse visual conditioning through spatial structure [Controlnet, t2i-adapter, uniControl], visual features [ipAdapter, anydoor, paintbyexample], and style transfer [instantstyle, stylecrafter, b-lora]. Most recently, large-scale DiT-based models like Flux Kontext [flux1kontext], Step1X-Edit [step1xedit], and Qwen-Image [qwenimage] employ architectures that separate text-stream and image-stream processing, using reference images as primary conditioning signals. Unlike T2I approaches requiring fine-tuning [dreambooth, Textualinversion, CustumDiff, perfectionnobody], these in-context image editing models directly condition generation on context images at inference time. While this offers powerful control and ease of use, it also introduces distinct privacy risks: personal images can be easily manipulated for identity impersonation or misinformation without owner consent, making protection mechanisms critically needed.
@@ 3.2 Context Propagates through Attention
Study I: Failure of the Standard Attack. First, we apply the standard PGD attack to our problem. We apply a standard untargeted attack (Eq. 3) to maximizing the flow-matching loss (Eq. 1) on Flux-Kontext. We term this baseline approach Diff-PGD (see Appendix).
@@ 5.1 Experimental Setup
Implementation Details. We set the step size to $\alpha=0.005$ and the noise budget to $\eta=0.1$, running a total of 800 optimization steps. At each training iteration, we randomly sample a diffusion timestep $t$ from the interval $\{980,\dots,1000\}$, and a prompt from a fixed pool of 60 prompts (complete list provided in the Appendix). Experiments are conducted on a single NVIDIA A800 80 GB GPU. Unless otherwise specified, these settings remain consistent across all experiments.
@@ Appendix B Extension Experiments on Items
Experiment Settings. We select 50 item images from the OminiControl Subject200K dataset, each resized to $\times512$ resolution. We follow the training settings used for human portraits and applied them to items, modifying only the prompt pool to better adapt them to item-related content. For the clean and perturbed version of each item, 10 outputs are generated per image-prompt pair with different seeds.

### v2

@@ 1 Introduction
However, the very strength of in-context personalization also introduces serious privacy risks. Since they operate at inference time and require only a single image, adversaries can easily exploit publicly shared photos to clone identities, mimic artistic styles, or reproduce proprietary content without consent. Consequently, as shown in Fig. LABEL:fig:results, the images users share online to express themselves may be weaponized to generate deepfakes, impersonation, and copyright violation, entirely bypassing conventional data protection mechanism.
@@ 2.1 Conditional Image Generation
Diffusion models have revolutionized controllable image synthesis [latentDM, SDXL, imagen, ediffi]. While early approaches rely on text prompts [GLIDE, DALLE2, cogview], recent advances enable diverse visual conditioning through spatial structure [Controlnet, t2i-adapter, uniControl], visual features [ipAdapter, anydoor, paintbyexample, ominicontrol, ominicontrol2], and style transfer [instantstyle, stylecrafter, b-lora]. Most recently, large-scale DiT-based models like Flux Kontext [flux1kontext], Step1X-Edit [step1xedit], and Qwen-Image [qwenimage] employ architectures that separate text-stream and image-stream processing, using reference images as primary conditioning signals. Unlike T2I approaches requiring fine-tuning [dreambooth, Textualinversion, CustumDiff, perfectionnobody], these in-context image editing models directly condition generation on context images at inference time. While this offers powerful control and ease of use, it also introduces distinct privacy risks: personal images can be easily manipulated for identity impersonation or misinformation without owner consent, making protection mechanisms critically needed.
@@ 3.2 Context Propagates through Attention
Study I: Failure of the Standard Attack. First, we apply the standard PGD attack to our problem. We apply a standard untargeted attack (Eq. 3) to maximizing the flow-matching loss (Eq. 1) on Flux-Kontext. We term this baseline approach Diff-PGD (see Appendix A.1).
@@ 5.1 Experimental Setup
Implementation Details. We set the step size to $\alpha=0.005$ and the noise budget to $\eta=0.1$, running a total of 800 optimization steps. At each training iteration, we randomly sample a diffusion timestep $t$ from the interval $\{980,\dots,1000\}$, and a prompt from a fixed pool of 60 prompts (complete list provided in the Appendix A.3). Experiments are conducted on a single NVIDIA A800 80 GB GPU. Unless otherwise specified, these settings remain consistent across all experiments.
@@ Appendix B Extension Experiments on Items
Experiment Settings. We select 50 item images from the OminiControl Subject200K dataset[ominicontrol], each resized to $\times512$ resolution. We follow the training settings used for human portraits and applied them to items, modifying only the prompt pool to better adapt them to item-related content. For the clean and perturbed version of each item, 10 outputs are generated per image-prompt pair with different seeds.